OpenStudy (anonymous):

(2p-q)2+q(4p-q) please write with work

OpenStudy (anonymous):

\[(2p-q)^2+q(4p-q)\] (2p-q)(2p-q) = 4p^2 -4pq +q^2 q(4p-q) = 4pq - q^2

OpenStudy (anonymous):

---> 4p^2 is the answer when you eliminate the rest.

OpenStudy (anonymous):

can you wite one by one

OpenStudy (anonymous):

Sure. \[(1) ~~~~~(2p-q)(2p-q) = 4p^2 -4pq +q^2\] \[(2)~~~~~ q(4p-q) = 4pq - q^2 \] \[(3)~~~ 4p^2-4pq+q^2+4pq-q^2\] \[(4)~~~~~4p^2 -4pq+4pq +q^2-q^2\] (5) cancel out the 4pq and q^2 and get the answer